OpenStudy (lilah96):

A manufacturing unit currently operates at 80 percent of its capacity. The profit function for the unit at the optimum output, x, is given by p(x) = -0.1x2 + 80x − 60. If the function f(x) models the current capacity of the unit, the composite function giving the unit's current profit function is f(p(x)) = -0.064x^2 + 6.4x - 60 p(f(x)) = -0.64x^2 + 0.64x - 60 p(f(x)) = -0.064x^2 + 64x - 60 p(f(x)) = -0.64x^2 + 6.4x - 60 f(p(x)) = -0.064x^2 + 64x - 60 . If the optimum output is 500 units, the current profit is $ 15,400 15,940 16,060 16,600 .

OpenStudy (lilah96):

f(p(x))=-0.064x^2 +64x-60 p(f(x))=0.64x^2 +6.4x-60 p(f(x))=-0.064X^ 64x-60 P(f(x))=0.64x^2 +0.64x-60 these are the chooses